# Redshift Assessment Infrastructure Layers (RAIL): Rubin-era photometric redshift stress-testing and at-scale production

### v2

@@ 5.3 Point-to-Point Metrics
  • PointOutlierRate computes the fraction of the $\Delta$ distribution outside of $[0,{\rm max}(0.06,3\sigma_{\rm iqr})]$. The upper limit is defined in the LSST Science Book (LSST Science Collaboration et al., 2009) and the lower bound of 0.06 is set to keep the fraction reasonable when $\sigma_{\rm IQR}$ is very small;
@@ 6.1 The Golden Spike: an end-to-end demonstration of RAIL
    Evaluating performance metrics: The Golden Spike notebook demonstrates RAIL’s ability to evaluate the performance metrics of the photo-$z$. We show the Probability Integral Transform function, a distribution-to-point metric described in Section 5.2, in Fig. 6. The PIT function (red line in the upper panel) is relatively close to the diagonal dashed line, which shows that FlexZBoost is providing photo-$z$ uncertainties that are consistent with the underlying uncertainties of the dataset.
@@ 6.3 Performance of RAIL
In this section, we briefly describe an initial study of the performance of RAIL. We note that the goal of this work is not to systematically benchmark the performance of the RAIL algorithms; we leave this to an ongoing comparative study of all RAIL algorithms (The Dark Energy Science Collaboration et al. in prep.).
@@
Algorithm Evaluation speed [$(s\,\times{\rm CPU})^{-1}$]
@@
Table 4: The speed in terms of [galaxies$(s\,\times{\rm CPU})^{-1}$] for four photo-$z$ algorithms. The speed is rounded to the second digit.
@@ 7 Summary
RAIL enables studies that address key challenges identified in DESC’s earlier photo-$z$ experiments, such as discrepancies between algorithms, inadequacies of traditional performance metrics, and the need for probabilistic approaches to model inherent redshift uncertainties (e.g., Moskowitz et al., 2024; Hang et al., 2024; Merz et al., 2024, Crawford et al. (in prep.)). Its is built around three types of modules: creation, which provides tools for generating mock photometric catalogs with tunable imperfections and realistic complexities; estimation, which supports a unified API for implementing and comparing a diverse array of algorithms to compute per-galaxy and ensemble photo-$z$ PDFs; and evaluation, which offers a flexible suite of metrics, including principled mathematical measures and science-case-specific performance evaluations.
@@
Future development will focus on expanding the range of supported algorithms, incorporating feedback from early users, and addressing emerging challenges in photo-$z$ systematics. Efforts will include refining methods for incorporating emerging algorithms, handling corner cases, and exploring integration with Rubin commissioning pipelines. For example, a major planned addition to RAIL is the SOMPZ method (Myles et al., 2021; Campos et al., 2024) adopted in the DES Y3 redshift calibration, which exploits the deep field photometry. This method also can be combined with clustering redshift (e.g., Giannini et al., 2024). Concerning the integration with the rest of the LSST analysis pipeline, efforts are currently being made to propagate the uncertainties associated with the photometry estimates of the methods described in this paper to cosmological constraints (Ruiz-Zapatero et al. In prep.). The aim of these efforts is to report photometric uncertainties in a way that can be ingested by DESC likelihood codes such as firecrown. This can be found in the DESC package nz_prior [^15] https://github.com/LSSTDESC/nz_prior.
@@ A.2 Tools
The photometry_tools include a few useful RAIL stages for photometry manipulation. Specifically, the HyperbolicMagnitudes allow the user to convert classical magnitudes and their respective errors to hyperbolic magnitudes (Lupton et al., 1999), given a smoothing parameter that is estimated in the stage HyperbolicSmoothing. The LSSTFluxToMagConverter converts the LSST fluxes and their respective errors into magnitudes. The Reddener and Dereddener compute (de)reddened magnitudes given a dust map, the RA and Dec of the catalog objects, as well as the wavelength-dependent extinction factors for each band, $A_{\lambda}/E(B-V)$. The default dust map is the SFD map (Schlegel et al., 1998; Schlafly & Finkbeiner, 2011) downloaded using the Python module dustmaps [^16] https://dustmaps.readthedocs.io/, but users can specify custom dust maps by specifying the paths to the maps.

### v3

@@ 5.3 Point-to-Point Metrics
  • PointOutlierRate computes the fraction of the $\Delta$ distribution outside of $0,{\rm max}(0.06,3\sigma_{\rm iqr})]$. The upper limit is defined in the LSST Science Book (LSST Science Collaboration et al., 2009) and the lower bound of 0.06 is set to keep the fraction reasonable when $\sigma_{\rm IQR}$ is very small;
@@ 6.1 The Golden Spike: an end-to-end demonstration of RAIL
    Evaluating performance metrics: The Golden Spike notebook demonstrates RAIL’s ability to evaluate the performance metrics of the photo-$z$. We show the Probability Integral Transform function, a distribution-to-point metric described in Section 5.2, in Fig. 6. The PIT function (red line in the upper panel) is relatively close to the diagonal dashed line, which shows that FlexZBoost is providing photo-$z$ uncertainties that are consistent with the underlying uncertainties of the dataset. Notice, however, that we do not set a ‘threshold’ value for a metric to be ‘good enough’. We leave this assessment to follow-up studies that forward-model these impacts to cosmological analyses.
@@ 6.3 Performance of RAIL
In this section, we briefly describe an initial study of the performance of RAIL. We note that the goal of this work is not to systematically benchmark the performance of the RAIL algorithms. There are ongoing efforts to compare the performance of all RAIL algorithms with realistic incomplete spectroscopic training sample using mock LSST data via the $p(z)$ and $n(z)$ data challenge (The Dark Energy Science Collaboration et al. in prep.). RAIL is applied in Rubin Data Preview 1 in SITCOMTN-154 (Charles et al., 2025) and Zhang et al. (2025). A systematic study of different metrics in the evaluation stage in RAIL is conducted by Crafford et al. in prep..
We benchmark the memory usage of selective RAIL algorithms, and summarize the maximum physical memory usage of each methods during training, estimation on 100,000 galaxies. We also record the model size of each method. All methods use 6 bands magnitude and magnitude error as input. The results are shown in Table 4. We note that the memory usage of kNN and LePhare are relatively large during training. We defer the memory optimization of these methods to future development.
Algorithm inform estimate model
BPz 0.177 0.401 0.0
CMNN 0.121 0.167 8.4
DNF 0.245 6.464 16.5
FlexZBoost 0.601 0.804 45.8
GPz 0.556 0.409 0.2
kNN 4.936 2.089 6.2
LePhare 11.842 1.078 413.7
simpleNN 0.292 0.213 0.0
TPz 0.462 0.858 87.1
Train-z 0.118 0.314 0.0
Table 4: Maximum physical memory (GB) and model size (MB) per algorithm and stage category.
@@
Algorithm Evaluation speed $(s\,\times{\rm CPU})^{-1}$]
@@
[…]
@@ 7 Summary
RAIL enables studies that address key challenges identified in DESC’s earlier photo-$z$ experiments, such as discrepancies between algorithms, inadequacies of traditional performance metrics, and the need for probabilistic approaches to model inherent redshift uncertainties (e.g., Moskowitz et al., 2024; Hang et al., 2024; Merz et al., 2024, Crawford et al. (in prep.)). It is built around three types of modules: creation, which provides tools for generating mock photometric catalogs with tunable imperfections and realistic complexities; estimation, which supports a unified API for implementing and comparing a diverse array of algorithms to compute per-galaxy and ensemble photo-$z$ PDFs; and evaluation, which offers a flexible suite of metrics, including principled mathematical measures and science-case-specific performance evaluations.
@@
Future development will focus on expanding the range of supported algorithms, incorporating feedback from early users, and addressing emerging challenges in photo-$z$ systematics. Efforts will include refining methods for incorporating emerging algorithms, handling corner cases, and exploring integration with Rubin commissioning pipelines. For example, a major planned addition to RAIL is the SOMPZ method (Myles et al., 2021; Campos et al., 2024) adopted in the DES Y3 redshift calibration, which uses a 2-tiered SOM to exploit deep field photometry and transfer the information to the wide field in a consistent, Bayesian approach. The resulting ensemble redshift distribution samples can also be combined with clustering redshift (e.g., Giannini et al., 2024) to yield a joint constraint on $n(z)$. Concerning the integration with the rest of the LSST analysis pipeline, efforts are currently being made to propagate the uncertainties associated with the photometry estimates of the methods to cosmological constraints. This is encompassed by the DESC package nz_prior [^15] https://github.com/LSSTDESC/nz_prior, which maps samples of $n(z)$ from RAIL to priors of chosen redshift uncertainty model, such as the ‘shift mean’ model, and fully connects RAIL with DESC likelihood codes such as firecrown. In a follow-up paper, Ruiz-Zapatero et al. (in prep.) aims to report impact of different photometric uncertainty models on cosmology using nz_prior. Finally, There are ongoing efforts to validate and quantitatively compare the performance of different algorithms within RAIL under realistic conditions, including scenarios with incomplete and non-representative training data. These efforts are primarily carried out through LSST DESC internal data challenges, which explicitly address algorithm performance at both the per-object level $p(z)$ and the ensemble level $n(z)$. The outcomes of these challenges will be presented in a series of follow-up papers.
@@
L. Medina-Varela: Memory profiling of RAIL algorithms.
@@ A.2 Tools
The photometry_tools include a few useful RAIL stages for photometry manipulation. Specifically, the HyperbolicMagnitudes allow the user to convert classical magnitudes and their respective errors to hyperbolic magnitudes (Lupton et al., 1999), given a smoothing parameter that is estimated in the stage HyperbolicSmoothing. The LSSTFluxToMagConverter converts the LSST fluxes and their respective errors into magnitudes. The Reddener and Dereddener compute (de)reddened magnitudes given a dust map, the RA and Dec of the catalog objects, as well as the wavelength-dependent extinction factors for each band, $A_{\lambda}/E(B-V)$. The default dust map is the SFD map (Schlegel et al., 1998; Schlafly and Finkbeiner, 2011) downloaded using the Python module dustmaps [^16] https://dustmaps.readthedocs.io/, but users can specify custom dust maps by specifying the paths to the maps.